Exercise 2. In a city, 65% of people drink coffee, 50% drink tea, and 25%
both. What is the probability that
(1) a person chosen at random will drink at least one of coffee or tea?
(2) a person will drink neither?
1
Expert's answer
2017-12-25T04:58:23-0500
Exercise 2:- Define events
A: person drinks coffee
B: person drinks tea
P(A) = 0.65, P(B) = 0.5, P(A and B) = 0.25

(1) P(A or B) = P(A) + P(B) - P(A and B)
= 0.65 + 0.5 - 0.25
= 0.9
(2) Required Probability is 1 - P(A or B). So, answer is 0.1